OpenStudy (anonymous):

Find the slope and y-intercept of the line. y = 2x + 5

OpenStudy (sheraz12345):

y=mx+c where m is the slope so what will be the slope .. Do u know.. U should know

OpenStudy (anonymous):

no i don't know :(

OpenStudy (the_fizicx99):

Y = mx + b mx is the slope, b is the y intercept so here: y = 5x + 7 5x is the slope 7 is my y intercept, get it?

OpenStudy (anonymous):

yea i get it so far

OpenStudy (sheraz12345):

@tHe_FiZiCx99 5x is not the slope only 5 is the slope

OpenStudy (anonymous):

?

OpenStudy (the_fizicx99):

Boi mx is the slope, shash

OpenStudy (sheraz12345):

You got it wrong bro @tHe_FiZiCx99 only m is the slope not mx .. X is the point. :)

OpenStudy (anonymous):

so could one of you please tell me the answer lol :)

OpenStudy (the_fizicx99):

Yesh it is ma bad

OpenStudy (sheraz12345):

No prob @tHe_FiZiCx99 ..@PeytonPerez372 2 compare ur question equation with y=mx+c and check what is m?

OpenStudy (anonymous):

m is 2???